OpenStudy (mattduhhelper):

On the coldest night in Vostok Station, Antarctica, the temperature dropped from 0°C to −24°C−24°C in 4 hours. What is the average temperature change per hour? Enter your answer in the box.

OpenStudy (mattduhhelper):

@T-Dawg02

OpenStudy (jiteshmeghwal9):

Average of any physical quantity is defined over an interval of time. so average temperature = \[\frac{\text{Final temperature - initial temperature}}{\text{Total time taken }}\]

OpenStudy (jiteshmeghwal9):

\[\frac{-24-0}{4}=?\]